# Minimum Bayes Risk Decoding for Error Span Detection in Reference-Free Automatic Machine Translation Evaluation

### v2

@@ Abstract
Error Span Detection (ESD) extends automatic machine translation (MT) evaluation by localizing translation errors and labeling their severity. Current generative ESD methods typically use Maximum a Posteriori (MAP) decoding, assuming that the model-estimated probabilities are perfectly correlated with similarity to the human annotation, but we often observe higher likelihood assigned to an incorrect annotation than to the human one.
We instead apply Minimum Bayes Risk (MBR) decoding to generative ESD. We use a sentence- or span-level similarity function for MBR decoding, which selects candidate hypotheses based on their approximate similarity to the human annotation. Experimental results on the WMT24 Metrics Shared Task show that MBR decoding significantly improves span-level performance and generally matches or outperforms MAP at the system and sentence levels. To reduce the computational cost of MBR decoding, we further distill its decisions into a model decoded via greedy search, removing the inference-time latency bottleneck.This work was done during the first author’s internship at National Institute of Information and Communications Technology, Kyoto, Japan. [^1] Our models and code are available at: https://github.com/vlaks425/MBR-ESD
@@ 1 Introduction
Automatic metrics for machine translation (MT) provide a fast proxy for costly human evaluation. From early n-gram metrics such as BLEU (Papineni et al., 2002) to learned metrics such as COMET (Rei et al., 2022), correlations with human judgments have improved at the system and sentence levels. However, most metrics still output a single scalar score per sentence (Freitag et al., 2024), hiding diagnostic information such as error location and severity that is crucial for model analysis and post-editing.
@@
State-of-the-art, LLM-based generative ESD methods (Freitag et al., 2024) typically rely on Maximum a Posteriori (MAP) decoding. The underlying assumption of MAP decoding is that the model-estimated probabilities are perfectly correlated with similarity to the human annotation. However, this assumption does not always hold. As illustrated in Table 1, we observe instances where the model log-likelihood is higher for an incorrect annotation produced by MAP decoding than for the correct human annotation, signifying a model error. This discrepancy, consistent with findings in other text generation tasks (Freitag et al., 2022a; Liu et al., 2022), leads us to question the optimality of MAP as a decoding objective for ESD.
Minimum Bayes Risk (MBR) decoding is a reranking method widely studied in MT (Kumar and Byrne, 2004; Tromble et al., 2008) and speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000). In this work, we apply MBR decoding to the ESD model with the objective of maximizing utility to the human annotation, independent of the model-estimated probabilities. Since the human annotation is unavailable at test time, we approximate it with a set of model-generated “support hypotheses” and select the hypothesis with the highest expected utility against this set.
We define the utility function as a similarity function over pairs of hypotheses and consider both sentence-level and span-level variants. The proposed sentence-level utility function measures similarity based on the difference between the severity-weighted error span counts of the two hypotheses. For the span-level utility function, we examine the official $\textsc{F1}(\cdot,\cdot)$ metric [^2] Referred to as the “meta-metric” in prior work. from the WMT QE Shared Task (Blain et al., 2023). However, we identify a defect in $\textsc{F1}(\cdot,\cdot)$ related to its handling of empty (error-free) annotations, which prevents it from performing soft evaluation. We therefore propose $\textsc{SoftF1}(\cdot,\cdot)$, a novel utility function that enables soft evaluation by addressing this defect.
@@ 2.2 Evaluation Metrics for ESD
Soft Pairwise Accuracy (SPA) (Thompson et al., 2024) and Pairwise Accuracy with Tie Calibration (${\text{Acc}}_{\text{eq}}^{*}$) (Deutsch et al., 2023) measure ranking consistency between model-predicted ESD annotations and the human annotation at the system and sentence levels, respectively.
At the span-level, to our knowledge, the only existing metric for ESD is the official metric of the WMT QE Shared Task (Blain et al., 2023), $\textsc{F1}(\cdot,\cdot)$, which computes character-level precision and recall between a hypothesis and a reference annotation. This metric can also be used as an MBR utility function when the reference is a support hypothesis.
However, we identify a critical defect in $\textsc{F1}(\cdot,\cdot)$, namely its handling of empty (error-free) annotations. As detailed in Appendix A.1, when the reference annotation is empty, $\textsc{F1}(\cdot,\cdot)$ returns its minimum value for any non-empty hypothesis, regardless of how minor the error might be. This is problematic for two reasons. First, as an evaluation metric, it fails to provide a soft penalty when the human annotation is error-free. Second, as an MBR utility function, it assigns minimum utility to any non-empty candidate whenever the support hypothesis is empty. Given the diversity of support hypotheses (Kamigaito et al., 2025) and the growing prevalence of perfect translations (Freitag et al., 2024), cases with empty human annotations or empty support hypotheses are likely to be frequent. We therefore regard this behavior as a significant limitation, which motivates our soft alternative, $\textsc{SoftF1}(\cdot,\cdot)$, introduced in Section 5.2.3.
@@ 2.3 MBR Decoding
MBR decoding has a long history in speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000), word alignment (Kumar and Byrne, 2002), and statistical MT (Kumar and Byrne, 2004; Tromble et al., 2008). Recently, several studies have revisited MBR decoding for text generation, demonstrating promising results (Wu et al., 2025; Heineman et al., 2024; Lyu et al., 2025). These studies suggest that MBR decoding can help overcome some of the limitations inherent to MAP. Furthermore, the properties (Müller and Sennrich, 2021; Kamigaito et al., 2025) and efficiency (Cheng and Vlachos, 2023; Vamvas and Sennrich, 2024; Deguchi et al., 2024) of MBR decoding remain active areas of research.
@@ 3.2 MBR Decoding
Unlike MAP, MBR explicitly optimizes task utility rather than model-estimated probability. Let $\mathcal{Y}$ denote the space of (unknown) ground-truth output $y$ for input $x$ and let $u(\cdot,\cdot)$ be a utility function, MBR decoding selects the hypothesis that minimizes the expected risk—or equivalently, maximizes the expected utility—with respect to $\mathcal{Y}$:
@@ 5.1 MBR Decoding for ESD
We specify the utility functions $u(\cdot,\cdot)$ in the next section.
@@ 5.2.1 MBR-ScoreSim
We introduce MBR-ScoreSim, which employs our proposed $\textsc{ScoreSim}(\cdot,\cdot)$ as the utility function. $\textsc{ScoreSim}(\cdot,\cdot)$ is a sentence-level utility function that measures the similarity between two hypotheses based on a severity-weighted error span count.
@@
Here $w_{\text{maj}}<0$ and $w_{\text{min}}<0$ are severity-dependent penalties, and $\alpha<0$ is a lower bound on the total penalty.
@@ 6 Experiments on MBR Decoding
Table 2: Evaluation results for the WMT 2024 Metrics Shared Task, conducted with Llama-3.3-70B-Inst and Gemma2-27b-it. We employ SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1 as metrics at the system, sentence, and span-levels, respectively. We report averaged scores across all translation directions. The best for the same model with NN is in bold. indicates significantly better performance than the MAP baseline with the same NN (# of candidate hypotheses) on all translation directions ($p<0.05$).
@@ 6.1 Experimental Setup
Following the WMT Metrics Shared Task, we used SPA (Thompson et al., 2024) and ${\text{Acc}}_{\text{eq}}^{*}$ (Deutsch et al., 2023) for system- and sentence-level evaluation, respectively. For the span level, we used $\textsc{SoftF1}(\cdot,E_{y})$ as the metric and reported the average scores over test instances. We selected $\textsc{SoftF1}(\cdot,E_{y})$ over $\textsc{F1}(\cdot,E_{y})$ as we consider it a more robust metric for ESD, particularly for handling empty annotations. For the SPA and ${\text{Acc}}_{\text{eq}}^{*}$, we employed the PERM-BOTH test (Deutsch et al., 2021). For SoftF1, [^4] SoftF1 indicates the average of the return values by $\textsc{SoftF1}(\cdot,E_{y})$ during evaluation across the entire dataset. Similar definitions apply to other functions. , since it operates at the instance level, we used paired bootstrap resampling (Koehn, 2004). We used MAP as the baseline for all statistical significance tests. For all ESD methods, sentence-level scores were derived using Eq. 6, system-level scores were calculated by averaging all sentence-level scores within that MT system.
@@ 6.2 Experimental Results
Drawing on the results presented in Table 2 and 3, we compare the performance of our proposed MBR instantiations against the MAP baseline. Our key findings are detailed below.
@@
At $N=256$, MBR-SoftF1 consistently improves SoftF1 over MAP for all four models, with these improvements being statistically significant in every setup; for SPA and ${\text{Acc}}_{\text{eq}}^{*}$, MBR-SoftF1 typically matches or outperforms MAP, although these gains are not always statistically significant. For example, with Mistral-Large-Inst, the SPA and ${\text{Acc}}_{\text{eq}}^{*}$ scores under MBR-SoftF1 are numerically close to those of MAP and the differences are not significant. At $N=16$, MBR-SoftF1 performed comparably to MAP with only marginal fluctuations, highlighting a key limitation: MBR-SoftF1 effectively trades computational cost for quality, requiring large sets of hypotheses ($N$). To address this resource dependency, we investigate MBR distillation in §8.
@@ 7.1 MBR Performance Upper Bound
Table 4: Understanding the performance upper bound of MBR through the oracle setup. Analysis is conducted under the sub-setup of Llama-3.3-70B-Inst with $N=256$. Model-* indicates MBR based on model-generated support hypotheses, while Oracle-* indicates MBR based on the human annotation.
@@ 7.2 MBR-SoftF1 Scalability
Table 5: Analysis of MBR-SoftF1 scalability. Analysis is conducted under the sub-setup of Llama-3.3-70B-Inst.
Due to computational constraints, we performed this analysis exclusively on the Llama-3.3-70B-Inst sub-setup, increasing $N$ to $1024$. As presented in Table 5, MBR-SoftF1 shows negligible performance gains at $N=1024$ compared to $N=256$ on SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1. We hypothesize that the model fails to generate sufficient diversity among candidates with the current temperature and Top-$K$ settings to benefit from the larger sample size. While tuning these hyperparameters for larger $N$ could potentially unlock further gains, we leave this investigation for future work.
@@ 8 Experiment on MBR Distillation
Despite the significant performance gains demonstrated in §6, MBR decoding comes at a high computational cost. Specifically, it relies on generating and scoring a large set of hypotheses (i.e., a large $N$), which incurs substantial latency and compute overhead during inference. To alleviate this bottleneck, we investigate MBR distillation, a technique designed to transfer the superior performance of MBR decoding into the model’s weights, thereby enabling comparable results using efficient greedy search.
@@ 8.1 Experimental Setup
The training dataset was derived from $(s,t)$ pairs in the WMT20–22 Metrics Shared Task datasets (Mathur et al., 2020; Freitag et al., 2021b; Freitag et al., 2022b). We used the WMT24 Metrics Shared Task as the test set. We used Llama-3.3-70B-Inst as the model. To generate the training targets, we employed the same MBR setup as in §6: $N=256$ with MBR-SoftF1. We refer to the model distilled via DPO as Distill-Greedy, which employs standard greedy decoding at inference time. Model training details are provided in Appendix C.

### v3

@@ Abstract
Error Span Detection (ESD) extends automatic machine translation (MT) evaluation by localizing translation errors and labeling their severity. Current generative ESD methods typically use Maximum a Posteriori (MAP) decoding, assuming that the model-estimated probabilities are perfectly correlated with similarity to the human annotation, but we often observe a higher likelihood assigned to an incorrect annotation than to the human one.
In contrast, we apply Minimum Bayes Risk (MBR) decoding to generative ESD. We use a sentence- or span-level similarity function for MBR decoding, which selects candidate hypotheses based on their approximate similarity to the human annotation. Experimental results on the WMT24 Metrics Shared Task show that MBR decoding significantly improves span-level performance and generally matches or outperforms MAP at the system and sentence levels. To reduce the computational cost of MBR decoding, we further distill its decisions into a model decoded via greedy search, removing the inference-time latency bottleneck.This work was done during the first author’s internship at National Institute of Information and Communications Technology, Kyoto, Japan. [^1] Our models and code are available at: https://github.com/vlaks425/MBR-ESD
@@ 1 Introduction
Automatic metrics for machine translation (MT) provide a fast proxy for costly human evaluation. From early n-gram metrics such as BLEU (Papineni et al., 2002) to learned metrics such as COMET (Rei et al., 2022), correlations with human judgments have improved at the system and sentence levels. However, most metrics still output a single scalar score per sentence (Freitag et al., 2024), hiding diagnostic information such as error location and severity, which are crucial for model analysis and post-editing.
@@
Current state-of-the-art LLM-based generative ESD methods (Freitag et al., 2024) typically rely on Maximum a Posteriori (MAP) decoding. The underlying assumption of MAP decoding is that the model-estimated probabilities are perfectly correlated with similarity to the human annotation. However, this assumption does not always hold. As illustrated in Table 1, we observe instances where the model log-likelihood is higher for an incorrect annotation produced by MAP decoding than for the correct human annotation, signifying a model error. This discrepancy, consistent with findings in other text generation tasks (Freitag et al., 2022a; Liu et al., 2022), leads us to question the optimality of MAP as a decoding objective for ESD.
Minimum Bayes Risk (MBR) decoding is a reranking method widely studied in MT (Kumar and Byrne, 2004; Tromble et al., 2008) and speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000). In this work, we apply MBR decoding to the ESD model with the objective of maximizing utility against the human annotation, independent of the model-estimated probabilities. Since the human annotation is unavailable at test time, we approximate it with a set of model-generated “support hypotheses” and select the hypothesis with the highest expected utility against this set.
We define the utility function as a similarity function over hypotheses pairs and consider both sentence-level and span-level variants. The proposed sentence-level utility function measures similarity based on the difference between the severity-weighted error span counts of the two hypotheses. For the span-level utility function, we examine the official $\textsc{F1}(\cdot,\cdot)$ metric (referred to as the “meta-metric” in prior work) from the WMT QE Shared Task (Blain et al., 2023). However, we identify a defect in $\textsc{F1}(\cdot,\cdot)$ related to its handling of empty (i.e., error-free) annotations, which prevents it from performing soft evaluation. We therefore propose $\textsc{SoftF1}(\cdot,\cdot)$, a novel utility function that enables soft evaluation by addressing this defect.
@@ 2.2 Evaluation Metrics for ESD
Soft Pairwise Accuracy (SPA) (Thompson et al., 2024) and Pairwise Accuracy with Tie Calibration (${\text{Acc}}_{\text{eq}}^{*}$) (Deutsch et al., 2023) measure ranking consistency. Specifically, SPA operates at the system level by evaluating pairwise rankings of different MT systems based on aggregated scores, whereas ${\text{Acc}}_{\text{eq}}^{*}$ functions at the sentence level, assessing the relative ordering of individual translation sentence pairs.
At the span level, to our knowledge, the only existing metric for ESD is the official metric of the WMT QE Shared Task (Blain et al., 2023), $\textsc{F1}(\cdot,\cdot)$, which computes character-level precision and recall between a hypothesis and a reference annotation. This metric can also be used as an MBR utility function when the reference is a support hypothesis.
However, we identify a critical defect in $\textsc{F1}(\cdot,\cdot)$, namely its handling of empty annotations. As detailed in Appendix A.1, when the reference annotation is empty, $\textsc{F1}(\cdot,\cdot)$ returns its minimum value for any non-empty hypothesis, regardless of how minor the error might be. This is problematic for two reasons. First, as an evaluation metric, it fails to provide a soft penalty when the human annotation is error-free. Second, as an MBR utility function, it assigns minimum utility to any non-empty candidate whenever the support hypothesis is empty. Given the diversity of support hypotheses (Kamigaito et al., 2025) and the growing prevalence of perfect translations (Freitag et al., 2024), cases with empty human annotations or empty support hypotheses are likely to be frequent. We therefore regard this behavior as a significant limitation, which motivates our soft alternative, $\textsc{SoftF1}(\cdot,\cdot)$, introduced in Section 5.2.3.
@@ 2.3 MBR Decoding
MBR decoding has a long history in speech recognition (Stolcke et al., 1997; Goel and Byrne, 2000), word alignment (Kumar and Byrne, 2002), and statistical MT (Kumar and Byrne, 2004; Tromble et al., 2008). Recently, several studies have revisited MBR decoding for text generation, demonstrating promising results (Wu et al., 2025; Heineman et al., 2024; Lyu et al., 2025). These studies suggest that MBR decoding can help overcome some of the limitations inherent in MAP. Furthermore, the properties (Müller and Sennrich, 2021; Kamigaito et al., 2025) and efficiency (Cheng and Vlachos, 2023; Vamvas and Sennrich, 2024; Deguchi et al., 2024) of MBR decoding remain active areas of research.
@@ 3.2 MBR Decoding
Unlike MAP, MBR explicitly optimizes task utility rather than model-estimated probability. Let $\mathcal{Y}$ denote the space of the (unknown) ground-truth output $y$ for input $x$ and $u(\cdot,\cdot)$ be a utility function, MBR decoding selects the hypothesis that minimizes the expected risk—or equivalently, maximizes the expected utility—with respect to $\mathcal{Y}$:
@@ 5.1 MBR Decoding for ESD
We specify the utility functions $u(\cdot,\cdot)$ below.
@@ 5.2.1 MBR-ScoreSim
We introduce MBR-ScoreSim, which employs our proposed $\textsc{ScoreSim}(\cdot,\cdot)$ as the utility function. $\textsc{ScoreSim}(\cdot,\cdot)$ is a sentence-level utility function that measures the similarity between two hypotheses based on severity-weighted error span counts.
@@
Here, $w_{\text{maj}}$ and $w_{\text{min}}$ are negative severity weights ($w<0$). These negative values reflect the subtractive nature of the MQM protocol (Freitag et al., 2021a), which penalizes the accumulation of error spans rather than rewarding quality. Finally, $\alpha<0$ defines a lower bound on the total penalty.
@@ 6 Experiments on MBR Decoding
Table 2: Evaluation results for the WMT 2024 Metrics Shared Task, conducted with Llama-3.3-70B-Inst and Gemma2-27b-it. We employ SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1 as metrics at the system, sentence, and span levels, respectively. We report average scores across all translation directions. The best score for the same model with NN is in bold. indicates significantly better performance than the MAP baseline with the same NN (# of candidate hypotheses) on all translation directions ($p<0.05$).
@@ 6.1 Experimental Setup
Following the WMT Metrics Shared Task, we used SPA (Thompson et al., 2024) and ${\text{Acc}}_{\text{eq}}^{*}$ (Deutsch et al., 2023) for system- and sentence-level evaluation, respectively. For the span level, we used $\textsc{SoftF1}(\cdot,E_{y})$ as the metric and reported the average scores over test instances. We selected $\textsc{SoftF1}(\cdot,E_{y})$ over $\textsc{F1}(\cdot,E_{y})$ as we consider it a more robust metric for ESD, particularly for handling empty annotations. For the SPA and ${\text{Acc}}_{\text{eq}}^{*}$, we employed the PERM-BOTH test (Deutsch et al., 2021). For SoftF1333SoftF1 indicates the average of the return values by $\textsc{SoftF1}(\cdot,E_{y})$ during evaluation across the entire dataset. Similar definitions apply to other functions., since it operates at the instance level, we used paired bootstrap resampling (Koehn, 2004). We used MAP as the baseline for all statistical significance tests. For all ESD methods, sentence-level scores were derived using Eq. 6, system-level scores were calculated by averaging all sentence-level scores within that MT system.
@@ 6.2 Experimental Results
Drawing on the results presented in Table 2 and 3, we compare the performance of our proposed MBR decoding instantiations against the MAP baseline. Our key findings are detailed below.
@@
At $N=256$, MBR-SoftF1 consistently improves SoftF1 over MAP, with these improvements being statistically significant in every setup; for SPA and ${\text{Acc}}_{\text{eq}}^{*}$, MBR-SoftF1 typically matches or outperforms MAP, although these gains are not always statistically significant. For example, with Mistral-Large-Inst, the SPA and ${\text{Acc}}_{\text{eq}}^{*}$ scores under MBR-SoftF1 are numerically close to those of MAP and the differences are not significant. Furthermore, at $N=16$, MBR-SoftF1 performed comparably to MAP with only marginal fluctuations, highlighting a key limitation: MBR-SoftF1 effectively trades computational cost for quality, requiring large sets of hypotheses ($N$). To address this resource dependency, we investigate MBR distillation in §8.
@@ 7.1 MBR Performance Upper Bound
Table 4: Performance upper bound of MBR identified through an oracle setup. Analysis is conducted using Llama-3.3-70B-Inst with $N=256$. Model-* indicates MBR based on model-generated support hypotheses, while Oracle-* indicates MBR based on the human annotation.
@@ 7.2 MBR-SoftF1 Scalability
Table 5: Analysis of MBR-SoftF1 scalability. Analysis is conducted using Llama-3.3-70B-Inst.
Due to computational constraints, we performed this analysis exclusively on the Llama-3.3-70B-Inst, increasing $N$ to $1024$. As presented in Table 5, MBR-SoftF1 shows negligible performance gains at $N=1024$ compared to $N=256$ on SPA, ${\text{Acc}}_{\text{eq}}^{*}$, and SoftF1. We hypothesize that the model fails to generate sufficient diversity among candidates with the current temperature and Top-$K$ settings to benefit from the larger sample size. While tuning these hyperparameters for larger $N$ could potentially unlock further gains, we leave this investigation for future work.
@@ 8 Experiment on MBR Distillation
Despite the significant performance gains demonstrated in §6, MBR decoding comes at a high computational cost. Specifically, it relies on generating and scoring a large set of hypotheses (i.e., a large $N$), which incurs substantial latency and compute overhead during inference. To alleviate this bottleneck, we investigate MBR distillation, a technique designed to transfer the superior performance of MBR decoding into the model weights, thereby enabling comparable results using efficient greedy search.
@@ 8.1 Experimental Setup
The training dataset was derived from $(s,t)$ pairs in the WMT20–22 Metrics Shared Task datasets (Mathur et al., 2020; Freitag et al., 2021b; Freitag et al., 2022b). We used the WMT24 Metrics Shared Task as the test set. We used Llama-3.3-70B-Inst as the model. To generate the training dataset, we employed the same MBR setup as in §6: $N=256$ with MBR-SoftF1. We refer to the model distilled via DPO as Distill-Greedy, which employs standard greedy decoding at inference time. See Appendix C for details.